# A Bombieri-Vinogradov theorem for sectors in real quadratic number fields

### v1

@@ 2.3. Vaughan’s identity
To treat sums of the form in (2.9) if $\mathcal{N}(\mathfrak{q})>\log^{A}x$, we use Vaughan’s identity for $\mathbb{K}$ to decompose the von Mangoldt function.
@@
$a_{3}(\mathfrak{n}):=-\sum\limits_{\begin{subarray}{c}\mathfrak{m}\mathfrak{r}=\mathfrak{n}\\ \mathcal{N}(\mathfrak{m})>U\\ \mathcal{N}(\mathfrak{r})>1\end{subarray}}\Lambda(\mathfrak{m})\sum\limits_{\begin{subarray}{c}\mathfrak{d}|\mathfrak{r}\\ \mathcal{N}(\mathfrak{d})\leq V\end{subarray}}\mu(\mathfrak{d}).$
@@ 2.4. Bessel’s inequality
Lemma 3 (Bessel’s inequality).
Let $\mathbf{v},\Phi_{1},\cdots,\Phi_{R}$ be vectors in an inner product space over the complex numbers. Then
[…]
where $||.||$ denotes the standard norm in the inner product space.
@@ 3. Reduction to sums with Hecke characters
3.1. Picking out residue class classes
@@ 3.2. Reduction to primitive characters
where the asterisk indicates that the relevant summation is restricted to primitive characters. We note that $\varphi(\mathfrak{j}\mathfrak{q}_{1})=\varphi(\mathfrak{j})\varphi(\mathfrak{q})$, which implies
@@ 3.3. Writing Dirichlet characters in terms of Hecke characters
(3.9) $\psi_{S}(y,\chi):=\sum\limits_{s\in S(\eta_{1},\eta_{2};y)}\Lambda(s)\chi_{\text{Hecke}}((s))\overline{\chi}_{\infty}(s),$
@@ 3.4. Picking out sectors
(3.15) $\sum\limits_{\begin{subarray}{c}\mathfrak{s}\in\mathcal{I}_{\mathbb{K}}\\ \mathcal{N}(\mathfrak{s})\leq y\end{subarray}}\Lambda(\mathfrak{s})(\chi_{\text{Hecke}}\xi_{\text{Hecke}}^{n})(\mathfrak{s}).$
@@
We now write $F(W)=F_{\chi}(W)$ for brevity and view $F(W)$ as a function in an indeterminate $W$. On the interval (1/2,1/2](-1/2,1/2], we may develop it into a Fourier series
(3.19) $F(W)=\sum\limits_{n\in\mathbb{Z}}a_{n}e(nW),$
@@
The equation (3.19) is valid for all $W\in(1/2,1/2]$ except for $W=\eta_{1}/2,\eta_{2}/2$, the discontinuities of the function $F(W)$. For $n\neq\tau$, we calculate that
@@
(3.20) $a_{n}\ll\frac{1}{1+|n|}\quad\mbox{for all }n\in\mathbb{N},$
where the implied constant is absolute. The following lemma provides a truncated version of the above Fourier series development.
@@
(3.22) $\psi_{S}(y,\chi)=\sum\limits_{|n|\leq Z}a_{n}(\chi)\sum\limits_{\begin{subarray}{c}\mathfrak{s}\in\mathcal{I}_{\mathbb{K}}\\ \mathcal{N}(\mathfrak{s})\leq y\end{subarray}}\Lambda(\mathfrak{s})(\chi_{\text{Hecke}}\xi_{\text{Hecke}}^{n})(\mathfrak{s})+O(\Sigma),$
@@ 3.5. Estimation of $\Sigma$
To estimate $\Sigma$, we need to obtain information about the spacing of $W(s)$, as $s$ runs over the sector S(1,1;x)S(-1,1;x). Our method will be an extension of that in [2, 4.2.4.] where $s$ was restricted to prime elements. We divide the sector S(1,1;x)S(-1,1;x), defined as in (1.3), into $O(\log x)$ dyadic parts
@@ 3.6. Application of the Siegel-Walfisz theorem
(3.33) $\begin{split}\sum\limits_{2\leq\mathcal{N}(\mathfrak{q})\leq Q}E^{\ast}(x;\mathfrak{q})\ll&(\log x)\sum\limits_{2\leq\mathcal{N}(\mathfrak{q})\leq Q}\frac{1}{\varphi(\mathfrak{q})}\;\;\sideset{}{{}^{\ast}}{\sum}\limits_{\chi\bmod\mathfrak{q}}\sum\limits_{|n|\leq x}\frac{1}{1+|n|}\times\\ &\max\limits_{y\leq x}\Bigg{|}\sum\limits_{\begin{subarray}{c}\mathfrak{s}\in\mathcal{I}_{\mathbb{K}}\\ \mathcal{N}(\mathfrak{s})\leq y\end{subarray}}\Lambda(\mathfrak{s})(\chi_{\text{Hecke}}\xi_{\text{Hecke}}^{n})(\mathfrak{s})\Bigg{|}+Qx^{1/2+2\varepsilon}.\end{split}$
Set $Q_{0}:=\log^{B}x$ where $B>0$ is arbitrarily given. Using Proposition 2, the contribution of the moduli $\mathfrak{q}$ with $2\leq\mathcal{N}(\mathfrak{q})\leq Q_{0}$ to the right-hand side of (3.33) is easily bounded by
(3.34) $(\log x)\sum\limits_{2\leq\mathcal{N}(\mathfrak{q})\leq Q_{0}}\frac{1}{\varphi(\mathfrak{q})}\;\;\sideset{}{{}^{\ast}}{\sum}\limits_{\chi\bmod\mathfrak{q}}\sum\limits_{|n|\leq Z}\frac{1}{1+|n|}\cdot\max\limits_{y\leq x}\Bigg{|}\sum\limits_{\begin{subarray}{c}\mathfrak{s}\in\mathcal{I}_{\mathbb{K}}\\ \mathcal{N}(\mathfrak{s})\leq y\end{subarray}}\Lambda(\mathfrak{s})(\chi_{\text{Hecke}}\xi_{\text{Hecke}}^{n})(\mathfrak{s})\Bigg{|}\ll_{A,B}\frac{x}{\log^{A}x}.$
Recall the assumption $Q\leq x^{2/3}$ from Theorem 1. We split the remaining contribution into $O(\log^{2}x)$ dyadic subsums $D(P,N)$ accounting for moduli $\mathfrak{q}$ satisfying $P<\mathcal{N}(\mathfrak{q})\leq 2P$ with $Q_{0}\leq P\leq Q$ and integers $n$ satisfying $N\leq|n|\leq 2N$ with $0\leq N\leq x$. If $Q\leq x$, then taking the well-known bound
@@
(3.35) $D(P,N)\ll\frac{\log^{2}x}{P(1+N)}\sum\limits_{\tilde{\chi}\in\mathcal{F}(P,N)}\max\limits_{y\leq x}\Big{|}\sum\limits_{\mathcal{N}(\mathfrak{s})\leq x}\Lambda(\mathfrak{s})\tilde{\chi}(\mathfrak{s})\Big{|},$
@@
Here it is understood that every Hecke character $\chi_{\text{Hecke}}$ belongs to precisely one primitive Dirichlet character modulo $\chi$. Following usual terminology, we refer to $T_{I}$ as a type I and $T_{II}$ as a type II bilinear sum. We note that
(3.37) $\sharp\mathcal{F}(P,N)\ll P^{2}(N+1)$
@@ 5. Large sieve inequality for Hecke characters in $\mathcal{F}(P,N)$
Let $\mathcal{F}(P,N)$ be defined as in (3.36), $K\geq 1$ and $a_{\mathfrak{k}}\in\mathbb{C}$ for any ideal $\mathfrak{k}\in\mathcal{I}_{\mathbb{K}}$. Then we have
(5.1) $\sum\limits_{\tilde{\chi}\in F(P,N)}\Big{|}\sum\limits_{\mathcal{N}(\mathfrak{k})\leq K}a_{\mathfrak{k}}\tilde{\chi}(\mathfrak{k})\Big{|}^{2}\ll\left(K+P^{8/3}\tilde{N}K^{1/3}\log^{2}2P\right)\sum\limits_{\mathcal{N}(\mathfrak{k})\leq K}|a_{\mathfrak{k}}|^{2}.$
@@
Let $\{\mathfrak{k}_{1},\cdots,\mathfrak{k}_{t}\}$ be the complete set of ideals $\mathfrak{k}\in\mathcal{I}_{\mathbb{K}}$ with norm less or equal $K$, and set
$\mathbf{v}:=(a_{\mathfrak{k}_{1}},...,a_{\mathfrak{k}_{t}}).$
Suppose that $\mathcal{F}(P,N)=\{\tilde{\chi}_{1},...,\tilde{\chi}_{R}\}$, and set
$\Phi_{r}:=(\tilde{\chi}_{r}(\mathfrak{k}_{1}),...,\tilde{\chi}_{r}(\mathfrak{k}_{t})).$
@@
$\sum\limits_{\tilde{\chi}\in F(P,N)}\Big{|}\sum\limits_{\mathcal{N}(\mathfrak{k})\leq K}a_{\mathfrak{k}}\tilde{\chi}(\mathfrak{k})\Big{|}^{2}=\sum\limits_{r\leq R}|(\mathbf{v},\Phi_{r})|^{2},$
and by Proposition 3 we have
[…]
If $\tilde{\chi}=\chi^{\prime}$, then trivially,
$\sum\limits_{\mathcal{N}(\mathfrak{k})\leq K}\tilde{\chi}\overline{\chi^{\prime}}(\mathfrak{k})\ll K.$
Now suppose that $\tilde{\chi}\not=\chi^{\prime}$. If $\tilde{\chi}$ is a Hecke character modulo $\mathfrak{q}$ and $\chi^{\prime}$ is a Hecke character modulo $\mathfrak{q}^{\prime}$, then $\tilde{\chi}\overline{\chi^{\prime}}$ is a non-principal Hecke character modulo $\mathfrak{q}\mathfrak{q}^{\prime}$. Therefore, Proposition 1 yields
$\sum\limits_{\mathcal{N}(\mathfrak{k})\leq K}\tilde{\chi}\overline{\chi^{\prime}}(\mathfrak{k})\ll P^{2/3}K^{1/3}\log^{2}2P.$
Combining everything above and recalling (3.37), the desired bound follows. ∎
Remark 3: We conjecture that (5.1) holds with the term $P^{8/3}\tilde{N}K^{1/3}\log^{2}2P$ replaced by $P^{2}\tilde{N}(PN)^{\varepsilon}$. This is because the size of the family $\mathcal{F}(P,N)$ of characters is $\asymp P^{2}\tilde{N}$.

### v2

@@ 2.3. Vaughan’s identity
To treat sums of the form in (2.9) if $\mathcal{N}(\mathfrak{q})>\log^{A}x$, we use Vaughan’s identity for $\mathcal{I}_{\mathbb{K}}$ to decompose the von Mangoldt function.
@@
$a_{3}(\mathfrak{n}):=-\sum\limits_{\begin{subarray}{c}\mathfrak{m}\mathfrak{r}=\mathfrak{n}\\ \mathcal{N}(\mathfrak{m})>U\\ \mathcal{N}(\mathfrak{r})>1\end{subarray}}\Lambda(\mathfrak{m})\sum\limits_{\begin{subarray}{c}\mathfrak{d}|\mathfrak{r}\\ \mathcal{N}(\mathfrak{d})\leq V\end{subarray}}\mu(\mathfrak{d}),$
with the sums above running over ideals in $\mathcal{I}_{\mathbb{K}}$.
@@ 2.4. Duality principle
Lemma 3 (Duality principle).
Let $\Phi:\mathbb{Z}^{2}\to\mathbb{C}$ be a complex-valued function. Suppose that for any sequence of complex numbers $(\beta_{n})_{n\leq N}$,
$\sum\limits_{m\leq M}\Big|\sum\limits_{n\leq N}\beta_{n}\Phi(m,n)\Big|^{2}\leq\Delta(M,N)\sum\limits_{n\leq N}|\beta_{n}|^{2}.$
Then for any sequence of complex numbers $\alpha_{m}$,
$\sum\limits_{n\leq N}\Big|\sum\limits_{m\leq M}\alpha_{m}\Phi(m,n)\Big|^{2}\leq\Delta(M,N)\sum\limits_{m\leq M}|\alpha_{m}|^{2},$
where $\Delta(M,N)$ is the same quantity in both inequalities.
@@ 3. Reduction to sums with Hecke characters
3.1. Picking out residue classes
@@ 3.2. Reduction to primitive characters
where the asterisk indicates that the relevant summation is restricted to primitive characters. We note that $\varphi(\mathfrak{j}\mathfrak{q}_{1})\geq\varphi(\mathfrak{j})\varphi(\mathfrak{q})$, which implies
@@ 3.3. Writing Dirichlet characters in terms of Hecke characters
(3.9) $\psi_{S}(y,\chi):=\sum\limits_{s\in S(\eta_{1},\eta_{2};y)}\Lambda(s)\chi_{\text{Hecke}}((s))\overline{\chi}_{\infty}(s).$
@@ 3.4. Picking out sectors
(3.15) $\sum\limits_{\begin{subarray}{c}\mathfrak{s}\in\mathcal{P}_{\mathbb{K}}\\ \mathcal{N}(\mathfrak{s})\leq y\end{subarray}}\Lambda(\mathfrak{s})(\chi_{\text{Hecke}}\xi_{\text{Hecke}}^{n})(\mathfrak{s}).$
@@
We now write $F(W)=F_{\chi}(W)$ for brevity and view $F(W)$ as a function in an indeterminate $W$. We note that $F(W)=0$ if $W\in(-1/2,1/2]\setminus(\eta_{1}/2,\eta_{2}/2]$. On the interval (1/2,1/2](-1/2,1/2], we may develop $F(W)$ into a Fourier series
$\sum\limits_{n\in\mathbb{Z}}a_{n}e(nW),$
@@
The equation
(3.19) $F(W)=\sum\limits_{n\in\mathbb{Z}}a_{n}e(nW)$
is valid for all $W\in(1/2,1/2]$ except for $W=\eta_{1}/2,\eta_{2}/2$, the discontinuities of the function $F(W)$. For $n\neq\tau$, we calculate that
@@
(3.20) $a_{n}\ll\frac{1}{1+|n|}\quad\mbox{for all }n\in\mathbb{Z},$
where the implied constant is absolute. The following lemma provides a truncated version of the above Fourier series expansion.
@@
(3.22) $\psi_{S}(y,\chi)=\sum\limits_{|n|\leq Z}a_{n}(\chi)\sum\limits_{\begin{subarray}{c}\mathfrak{s}\in\mathcal{P}_{\mathbb{K}}\\ \mathcal{N}(\mathfrak{s})\leq y\end{subarray}}\Lambda(\mathfrak{s})(\chi_{\text{Hecke}}\xi_{\text{Hecke}}^{n})(\mathfrak{s})+O(\Sigma),$
@@ 3.5. Estimation of $\Sigma$
To estimate $\Sigma$, we need to obtain information about the spacing of $W(s)$, as $s$ runs over the sector S(1,1;x)S(-1,1;x). Our method will be an extension of that in [2, subsection 4.2.4.] where $s$ was restricted to prime elements. We divide the sector S(1,1;x)S(-1,1;x), defined as in (1.3), into $O(\log x)$ dyadic parts
@@ 3.6. Application of the Siegel-Walfisz theorem
(3.33) $\begin{split}\sum\limits_{2\leq\mathcal{N}(\mathfrak{q})\leq Q}E^{\ast}(x;\mathfrak{q})\ll&(\log x)\sum\limits_{2\leq\mathcal{N}(\mathfrak{q})\leq Q}\frac{1}{\varphi(\mathfrak{q})}\;\;\sideset{}{{}^{\ast}}{\sum}\limits_{\chi\bmod\mathfrak{q}}\ \sum\limits_{|n|\leq x}\frac{1}{1+|n|}\times\\ &\max\limits_{y\leq x}\Bigg|\sum\limits_{\begin{subarray}{c}\mathfrak{s}\in\mathcal{P}_{\mathbb{K}}\\ \mathcal{N}(\mathfrak{s})\leq y\end{subarray}}\Lambda(\mathfrak{s})(\chi_{\text{Hecke}}\xi_{\text{Hecke}}^{n})(\mathfrak{s})\Bigg|+Qx^{1/2+2\varepsilon}.\end{split}$
Set $Q_{0}:=\log^{B}x$ where $B>0$ is arbitrarily given. If $2\leq\mathcal{N}(\mathfrak{q})\leq Q_{0}$, then we write
(3.34) $\sum\limits_{\begin{subarray}{c}\mathfrak{s}\in\mathcal{P}_{\mathbb{K}}\\ \mathcal{N}(\mathfrak{s})\leq y\end{subarray}}\Lambda(\mathfrak{s})(\chi_{\text{Hecke}}\xi_{\text{Hecke}}^{n})(\mathfrak{s})=\frac{1}{h}\cdot\sum\limits_{\psi\in\mathcal{X}}\sum\limits_{\begin{subarray}{c}\mathfrak{s}\in\mathcal{I}_{\mathbb{K}}\\ \mathcal{N}(\mathfrak{s})\leq y\end{subarray}}\Lambda(\mathfrak{s})(\chi_{\text{Hecke}}\xi_{\text{Hecke}}^{n}\psi)(\mathfrak{s}),$
where $\mathcal{X}$ is the group of class group characters, and use Proposition 2 to bound the character sum on the right-hand side. In this way, the contribution of the moduli $\mathfrak{q}$ with $2\leq\mathcal{N}(\mathfrak{q})\leq Q_{0}$ to the right-hand side of (3.33) is easily bounded by
(3.35) $(\log x)\sum\limits_{2\leq\mathcal{N}(\mathfrak{q})\leq Q_{0}}\frac{1}{\varphi(\mathfrak{q})}\;\;\sideset{}{{}^{\ast}}{\sum}\limits_{\chi\bmod\mathfrak{q}}\sum\limits_{|n|\leq Z}\frac{1}{1+|n|}\cdot\max\limits_{y\leq x}\Bigg|\sum\limits_{\begin{subarray}{c}\mathfrak{s}\in\mathcal{P}_{\mathbb{K}}\\ \mathcal{N}(\mathfrak{s})\leq y\end{subarray}}\Lambda(\mathfrak{s})(\chi_{\text{Hecke}}\xi_{\text{Hecke}}^{n})(\mathfrak{s})\Bigg|\ll_{A,B}\frac{x}{\log^{A}x}.$
Recall the assumption $Q\leq x^{4/5}$ from Theorem 1. We split the remaining contribution into $O(\log^{2}x)$ dyadic subsums $D(P,N)$ accounting for moduli $\mathfrak{q}$ satisfying $P<\mathcal{N}(\mathfrak{q})\leq 2P$ with $Q_{0}\leq P\leq Q$ and integers $n$ satisfying $N\leq|n|\leq 2N$ with $0\leq N\leq x$. If $Q\leq x$, then using (3.34) and taking the well-known bound
@@
(3.36) $D(P,N)\ll\frac{\log^{2}x}{P(1+N)}\sum\limits_{\tilde{\chi}\in\mathcal{F}(P,N)}\max\limits_{y\leq x}\Big|\sum\limits_{\begin{subarray}{c}\mathfrak{s}\in\mathcal{I}_{\mathbb{K}}\\ \mathcal{N}(\mathfrak{s})\leq y\end{subarray}}\Lambda(\mathfrak{s})\tilde{\chi}(\mathfrak{s})\Big|,$
@@
Here it is understood that every Hecke character $\chi_{\text{Hecke}}$ belongs to precisely one primitive Dirichlet character $\chi$. Following usual terminology, we refer to $T_{I}$ as a type I and $T_{II}$ as a type II bilinear sum. We note that
(3.38) $\sharp\mathcal{F}(P,N)\ll P^{2}(N+1).$
@@ 5. Large sieve inequality for Hecke characters in $\mathcal{F}(P,N)$
Let $\mathcal{F}(P,N)$ be defined as in (3.37), $K\geq 1$ and $a_{\mathfrak{k}}\in\mathbb{C}$ for any ideal $\mathfrak{k}\in\mathcal{P}_{\mathbb{K}}$. Then we have
(5.1) $\sum\limits_{\tilde{\chi}\in F(P,N)}\Big|\sum\limits_{\mathcal{N}(\mathfrak{k})\leq K}a_{\mathfrak{k}}\tilde{\chi}(\mathfrak{k})\Big|^{2}\ll\left(K+P^{3}\tilde{N}(PK)^{\varepsilon}\right)\sum\limits_{\mathcal{N}(\mathfrak{k})\leq K}|a_{\mathfrak{k}}|^{2}.$
@@
[…]
(5.2) $S:=\sum\limits_{\mathcal{N}(\mathfrak{k})\leq K}\Big|\sum\limits_{\tilde{\chi}\in F(P,N)}b_{\tilde{\chi}}\tilde{\chi}(\mathfrak{k})\Big|^{2}\ll\big(K+P^{3}\tilde{N}(PK)^{\varepsilon}\big)\,\sum\limits_{\tilde{\chi}\in F(P,N)}|b_{\tilde{\chi}}|^{2}.$
Set
$\omega(x):=\exp\left(-\frac{x^{2}}{K^{2}}\right).$
@@
$S\ll\sum\limits_{\mathfrak{k}\in\mathcal{I}_{\mathbb{K}}}\omega\left(\mathcal{N}(\mathfrak{k})\right)\cdot\Big|\sum\limits_{\tilde{\chi}\in F(P,N)}b_{\tilde{\chi}}\tilde{\chi}(\mathfrak{k})\Big|^{2}.$
Opening the square and interchanging the order of summations, we deduce that
(5.3) $\begin{split}S\ll&\sum\limits_{\tilde{\chi}_{1},\tilde{\chi}_{2}\in F(P,N)}b_{\tilde{\chi}_{1}}\overline{b_{\tilde{\chi}}}_{2}\sum\limits_{\mathfrak{k}\in\mathcal{I}_{\mathbb{K}}}\omega\left(\mathcal{N}(\mathfrak{k})\right)\tilde{\chi}_{1}\overline{\tilde{\chi}}_{2}(\mathfrak{k})\\ \ll&\sum\limits_{\tilde{\chi}_{1},\tilde{\chi}_{2}\in F(P,N)}\big(|b_{\tilde{\chi}_{1}}|^{2}+|b_{\tilde{\chi}_{2}}|^{2}\big)\Big|\sum\limits_{\mathfrak{k}\in\mathcal{I}_{\mathbb{K}}}\omega\left(\mathcal{N}(\mathfrak{k})\right)\tilde{\chi}_{1}\overline{\tilde{\chi}}_{2}(\mathfrak{k})\Big|\\ \ll&\sum\limits_{\tilde{\chi}_{1}\in F(P,N)}|b_{\tilde{\chi}_{1}}|^{2}\sum\limits_{\tilde{\chi}_{2}\in F(P,N)}\Big|\sum\limits_{\mathfrak{k}\in\mathcal{I}_{\mathbb{K}}}\omega(\mathcal{N}(\mathfrak{k}))\tilde{\chi}_{1}\overline{\tilde{\chi}}_{2}(\mathfrak{k})\Big|\\ \ll&\Big(\sum\limits_{\tilde{\chi}_{1}\in F(P,N)}|b_{\tilde{\chi}_{1}}|^{2}\Big)\max_{\tilde{\chi_{1}}\in F(P,N)}\sum\limits_{\tilde{\chi}_{2}\in F(P,N)}\Big|\sum\limits_{\mathfrak{k}\in\mathcal{I}_{\mathbb{K}}}\omega(\mathcal{N}(\mathfrak{k}))\tilde{\chi}_{1}\overline{\tilde{\chi}}_{2}(\mathfrak{k})\Big|.\end{split}$
Clearly,
(5.4) $\Big|\sum\limits_{\mathfrak{k}\in\mathcal{I}_{\mathbb{K}}}\omega(\mathcal{N}(\mathfrak{k}))\tilde{\chi}_{1}\overline{\tilde{\chi}}_{2}(\mathfrak{k})\Big|\leq\sum\limits_{\mathfrak{k}\in\mathcal{I}_{\mathbb{K}}}\omega(\mathcal{N}(\mathfrak{k}))\ll K\quad\mbox{if }\tilde{\chi}_{1}=\tilde{\chi}_{2}.$
If $\tilde{\chi}_{1}\not=\tilde{\chi}_{2}$, then applying the inverse Mellin transform, we may transform the above sum over $\mathfrak{k}$ into
(5.5) $\begin{split}\sum\limits_{\mathfrak{k}\in\mathcal{I}_{\mathbb{K}}}\omega(\mathcal{N}(\mathfrak{k}))\tilde{\chi}_{1}\overline{\tilde{\chi}}_{2}(\mathfrak{k})&=\frac{1}{2\pi i}\int\limits_{\Re(s)=\sigma}\tilde{\omega}(s)\sum\limits_{\mathfrak{k}\in\mathcal{I}_{\mathbb{K}}}\frac{\tilde{\chi}_{1}\overline{\tilde{\chi}}_{2}(\mathfrak{k})}{\mathcal{N}(\mathfrak{k})^{s}}ds\\ &=\frac{1}{2\pi i}\int\limits_{\Re(s)=\sigma}\tilde{\omega}(s)L(s,\tilde{\chi}_{1}\overline{\tilde{\chi}}_{2})ds,\end{split}$
where $\sigma$ is any real number greater than 1 and
$\tilde{\omega}(s):=\int\limits_{0}^{\infty}\omega(y)y^{s-1}dy=\frac{1}{2}\cdot K^{s}\Gamma\left(\frac{s}{2}\right)$
is the Mellin transform of $\omega$. Since both functions $\tilde{\omega}(s)$ and $L(s,\tilde{\chi}_{1}\overline{\tilde{\chi}}_{2})$ are analytic in the half plane $\Re(s)>0$, the line of integration can be shifted to $\Re(s)=\varepsilon>0$. By the rapid decay of the Gamma function on vertical lines, we have
(5.6) $\widetilde{\omega}(\varepsilon+it)\ll K^{\varepsilon}(1+|t|)^{-j}$
for any $j>0$. The convexity bound for Hecke $L$-functions (see [11, Exercise 3 on page 100]) implies that
(5.7) $L(\varepsilon+it,\tilde{\chi}_{1}\overline{\tilde{\chi}}_{2})\ll{\bf f}(\tilde{\chi}_{1}\overline{\tilde{\chi}}_{2})^{1/2+\varepsilon}(1+|t|)^{1+\varepsilon},$
where ${\bf f}(\tilde{\chi}_{1}\overline{\tilde{\chi}}_{2})$ is the norm of the conductor of $\tilde{\chi}_{1}\overline{\tilde{\chi}}_{2}$ (for information on the conductor, see [14, section 3.3]). Recalling (3.37), we have
(5.8) ${\bf f}(\tilde{\chi}_{1}\overline{\tilde{\chi}}_{2})\ll P^{2}$
if $\tilde{\chi}_{1},\tilde{\chi}_{2}\in F(P,N)$. Using (5.5), (5.6), (5.7) and (5.8) with $j>2+\varepsilon$, we have the bound
(5.9) $\sum\limits_{\mathfrak{k}\in\mathcal{I}_{\mathbb{K}}}\omega(\mathfrak{k})\tilde{\chi}_{1}\overline{\tilde{\chi}}_{2}(\mathfrak{k})\ll P(PK)^{\varepsilon}\quad\mbox{if }\tilde{\chi}_{1}\not=\tilde{\chi}_{2}.$
Combining (3.38), (5.3), (5.4) and (5.9), we obtain the desired estimate (5.2). This completes the proof. ∎
Remark 3: Since $\sharp\mathcal{F}(P,N)\ll P^{2}\tilde{N}$, we conjecture that (5.1) holds with the term $P^{3}\tilde{N}(PK)^{\varepsilon}$ replaced by $P^{2}\tilde{N}(P\tilde{N}K)^{\varepsilon}$.